your brother starts out 10 minutes before you on a bike path and rides at a rate of 9mph. if you ride at a rate of 12mph, how long will it take you to catch up to him
Answer by richwmiller(17219)

Brother got a 10 minute head start
In those ten minutes he went 9*1/6=9/6=1.5 miles
We are gaining on him at 3 mph.
So it will take 30 minutes to catch up
another way to do this
9*(x+1/6)=12*x
9x+1.5=12x
1.5=3x
1.5/3=x
1/2=x
x= 1/2 hr.